### v3

# Universal Approximation Theory: The Basic Theory for Transformer-based Large Language Models

Abstract
Language models have become a focal point in the field of artificial intelligence, especially following the emergence of ChatGPT. Large-scale Transformer networks have quickly become the primary method for advancing natural language processing algorithms. These models, built on the Transformer architecture, are capable of engaging in near-human-like interactions and can even assist in guiding human tasks. Despite their impressive capabilities, there remain some critical theoretical questions in the development of large language models (LLMs): What makes the Transformer architecture so effective in driving intelligent language applications, such as translation and coding? What underlies the in-context learning (ICL) capability of LLMs? How does the LoRA approach enhance the fine-tuning of LLMs? What supports the feasibility of pruning in LLMs? To address these key questions and explore the technical strategies within LLMs, we draw on the Universal Approximation Theory (UAT) to provide a theoretical foundation, revealing the mechanisms behind these advancements.
1 Introduction
In recent years, the rapid emergence of LLMs in the field of artificial intelligence has undoubtedly become one of the most notable advancements within the domain. The core allure of these models stems from their extraordinary capabilities in language processing. Language, as a unique crystallization of human intelligence, serves not only as the external reflection of thought but also as the bridge for communication, the cornerstone for the dissemination of knowledge, and the continuation of civilization, profoundly shaping the identity of humans as a unique species. Thus, endowing machines with the ability to understand and generate language marks a significant leap towards the realization of true artificial intelligence. The emergence of models such as the ChatGPT (Radford and Narasimhan 2018; Brown et al. 2020; Achiam et al. 2023), the Llama (Touvron et al. 2023), and the PaLM (Chowdhery et al. 2023) vividly demonstrates this point.
A distinctive feature of LLMs is their immense parameter size (Achiam et al. 2023; Touvron et al. 2023; Chowdhery et al. 2023; Chen et al. 2021; Zeng et al. 2022), often amounting to hundreds of billions or even trillions (for instance, GPT-3’s (Brown et al. 2020) 175 billion parameters and PaLM’s (Chowdhery et al. 2023) 540 billion parameters). This vast parameter scale lays the foundation for their exceptional language processing capabilities and enables them to exhibit almost human-like traits, such as ICL  (Brown et al. 2020; Dong et al. 2022), instruction following (Sanh et al. 2021; Ouyang et al. 2022; Wei et al. 2021), and multi-step reasoning (Wei et al. 2022). Notably, these colossal models are predominantly trained by tech giants like Google and Microsoft using large-scale GPU clusters (Zhao et al. 2023), sparking a research fervor on how to efficiently fine-tune them with limited GPU resources. The advent of Lora (Hu et al. 2021) fine-tuning technology has provided an effective pathway for this, allowing for the fine-tuning of large models under resource constraints without the need to adjust all parameters of the original model comprehensively. Moreover, model pruning techniques (Sun et al. 2023; Ma, Fang, and Wang 2023) are crucial for deploying large models in resource-constrained environments, aiming to reduce the model size for operation on smaller devices. Today, LLMs possess a diverse range of functionalities, from translation and text summarization to automatic code generation, demonstrating their versatility.
Despite the rapid advancements of LLMs towards higher intelligence and reliability, the theoretical foundations behind these models remain largely unexplored and shrouded in uncertainty. The scientific community is actively seeking to uncover the intrinsic mechanisms behind their powerful capabilities, including the analysis of ICL mechanisms (Xie et al. 2021; Min et al. 2022). In this context, we establish the UAT (Cybenko 2007; Hornik, Stinchcombe, and White 1989) as the mathematical essence of Transformer-based LLMs and use the perspective of UAT to elucidate key technologies and phenomena within LLMs, aiming to provide explanations from a new angle.
While previous studies have explored the relationship between Transformers and UAT (Yun et al. 2019, 2020; Kratsios et al. 2021; Alberti et al. 2023), these studies are based on a series of complex assumptions and derivations, lacking generality and failing to intuitively address some of the issues present in LLMs. In contrast, this paper aims to provide proof that bridges Transformers and UAT, grounded in the foundational mathematical form of UAT as established by Cybenko; Hornik, Stinchcombe, and White. Our proof is characterized by its simplicity, absence of assumptions, and generality, making it applicable to Convolutional Neural Networks (CNNs) as well (Wang and Li 2024b). Our approach not only provides intuitive answers to some of the challenges within LLMs but also aids in the design of deep learning networks from a theoretical standpoint (Wang and Li 2024a). These capabilities are not addressed by previous theories. Our contributions are as follows:
  • We prove that the Transformer is the tangible embodiment of UAT.
  • We deliver a rigorous scientific explanation of Transformer-based LLMs through the UAT.
  • We explain the characteristics of LLMs, such as ICL, instruction following, multi-step reasoning, and the technologies applied within LLMs like Lora, pruning, and LLMs’ strong generalization capabilities.
Our article is structured as follows: In Section 2, we begin by introducing the UAT and propose that to demonstrate the Transformer’s adherence to UAT, it is necessary to show that both Linear and Multi-Head Attention (MHA) can be represented in the form of matrix-vector multiplication. In Section 3, we establish that the Transformer belongs to the category of UAT. Our proof strategy unfolds starting with Section 3.1, where we present the idea to express Linear and MHA as matrix-vector products. In Section 3.2, we detail the computational methods required for the proof. Sections 3.3 and 3.4 are dedicated to proving that both Linear and MHA can indeed be represented as matrix-vector operations. Further in Section 4, we leverage UAT to theoretically elucidate some fundamental issues (Generalization 4.1, ICL 4.2) and techniques (Pruning 4.3 and LoRA 4.4) associated with LLMs. Finally, we conclude with a summary of the existing challenges and potential future developments for LLMs in Section 4.5.
2 The Universal Approximation Theory
Up to now, the UAT (Cybenko 2007) is the most widely recognized fundamental theory in deep learning. However, this theorem is only applicable to the simplest form of neural networks, the multilayer perceptron (MLP) (Cybenko 2007; Popescu et al. 2009). Due to the increased complexity of Transformer networks, they cannot be mathematically expressed in the same form as UAT. Thus, the theory has not yet been extended to Transformer networks. Our aim in this paper is to unify Transformer networks under the framework of UAT, thereby standardising their mathematical representation. Before unifying their mathematical forms, we provide a brief overview of the theorem, which was initially proposed by Cybenko (2007). This theorem encompasses numerous conclusions and proof details. Although it has been further developed, its fundamental mathematical form remains unchanged. Therefore, this paper explains the theory based on the UAT form presented by Cybenko (2007). Theorem 2 from Cybenko (2007) states that if $\sigma$ is any continuous sigmoidal function, then finite sums of the following form:
$G(\mathbf{x})=\sum_{j=1}^{N}\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)$ (1)
is dense in $C\left(\mathbf{I}_{n}\right)$. Here, $\mathbf{W}_{j}\in\mathbb{R}^{n}$ and $\alpha_{j},\theta\in\mathbb{R}$ are fixed. For any $f\in C\left(\mathbf{I}_{n}\right)$ and $\varepsilon>0$, there exists a function $G(\mathbf{x})$:
$|G(\mathbf{x})-f(\mathbf{x})|<\varepsilon\quad\text{ for all }\quad\mathbf{x}\in\mathbf{I}_{n}.$ (2)
This implies that, when $N$ is sufficiently large, a neural network can approximate any continuous function on a closed interval. Hornik, Stinchcombe, and White (1989) further demonstrates that multilayer feedforward networks also conform to the UAT, capable of approximating arbitrary Borel measurable functions. Observing Equation (1), where the function $G(\mathbf{x})$ yields a scalar output in $\mathbb{R}$, the scenario expands naturally when $G(\mathbf{x})$ maps to $\mathbb{R}^{m}$, requiring the approximation in each dimension. It becomes evident that to accommodate this multidimensional output, a simple extension to Equation (1) suffices: the transformation matrix $\mathbf{W}_{j}$ is revised to the space $\mathbb{R}^{n\times m}$, the bias term $\theta_{j}$ is recast as a vector in $\mathbb{R}^{m}$, and $\alpha_{j}$ is reshaped into a matrix. Nevertheless, in these formulas, the theorem does not straightforwardly apply to Transformer architectures. However, if we succeed in reformulating both the Linear ($\mathbf{x}|\mathbf{W}$) and MHA ($\mathbf{x}|\mathbf{W}$) components into a unified representation, $\mathbf{W}^{\prime}\mathbf{x}^{\prime}$, where $\mathbf{W}$ and $\mathbf{x}$ represent the parameters and inputs for each component respectively, and $\mathbf{W}^{\prime}$ is derived from $\mathbf{W}$ while $\mathbf{x}^{\prime}$ signifies a column vector derived from $\mathbf{x}$ (for details, refer to Section 3.1). It is easy to prove that multi-layer Transformer networks are concrete embodiments of UAT.
Consequently, our objective is to represent Linear and MHA architectures as a matrix-vector multiplication similar to Eq. (3). This approach offers two key benefits: Firstly, it allows for a fundamental mathematical investigation of deep learning models. Secondly, it provides a more intuitive and objective basis for comparing differences among various models. On this foundation, we propose to transform Linear and MHA into the matrix-vector format.
3 UAT for Transformer
In the previous section, we have outlined our fundamental goal to unify the Transformer model under the framework of UAT by converting Linear and MHA operations in the Transformer into matrix-vector forms. In this section, we will use the Matrix-Vector method to transform Transformer operations into matrix times vector forms.
3.1 Matrix-Vector Method
Figure 1: The transformation process of the Matrix-Vector Method.
Before delving into transforming Linear and MHA into their matrix-vector format, we introduce the Matrix-Vector Method, which will subsequently be employed to cast both Linear and MHA operations into a unified matrix-vector format. This method constitutes a strategic realignment of input data and corresponding parameters of various transformations within the network, as illustrated in Figure 1. The underlying principle is as follows: both the input ($\mathbf{x}$) and output ($\mathbf{y}$) data are reconfigured uniformly through a transformation $\mathbf{T_{D}}$ into column vectors ($\mathbf{x}^{\prime}$ and $\mathbf{y}^{\prime}$, while parameter tensor ($\mathbf{W}$) is reorganized into matrix form ($\mathbf{W}^{\prime}$). A critical requirement is that identical operations throughout the LLMs network adhere to the same restructuring scheme, thereby eliminating the need for additional transformations on intermediate feature data; they can directly be represented as column vectors. Additionally, by default, matrix variables in the original formulas are represented in bold, such as $\mathbf{x}$, and in the Matrix-Vector form, corresponding variables are denoted with a prime symbol () in the upper right corner, such as $\mathbf{x}^{\prime}$. Elements within matrices are represented by corresponding lowercase letters with subscripts, for example, $x_{i}$. The Matrix-Vector Method can be succinctly encapsulated as follows:
$\mathbf{y}=T(\mathbf{x}|\mathbf{W})\rightarrow\mathbf{y}^{\prime}=\mathbf{W}^{\prime}\mathbf{x}^{\prime}$ (3)
It is obvious that $T_{D}$ and $T_{P}$ are not fixed, we could design various kinds of ways to do those. For convenience, we present a methodology tailored for Linear in Section 3.3 and an approach for MHA in Section 3.4, thereby illustrating the adaptability and application of the Matrix-Vector Method across different components of the Transformer architecture.
3.2 Diamond Matrix Multiplication
While our goal is to transform Linear and MHA modules into a matrix-vector format, this process is complex. In order to describe this transformation process clearly, we propose a new computational method called the Diamond Multiplication Method, denoted by the symbol $\diamond$. (The purpose of proposing this method is to clearly illustrate the calculation process in turning MHA into Matrix-Vector format. If we use direct matrix multiplication, some computational details cannot be visually distinguished.)
$\left(\begin{array}[]{cccc}w_{1,1}&w_{1,2}&\cdots&w_{1,n}\\ w_{2,1}&w_{2,2}&&w_{2,n}\\ \vdots&\vdots&\vdots&\vdots\\ w_{m,1}&w_{m,2}&&w_{m,n}\end{array}\right)\diamond\left(\begin{array}[]{c}x_{1}\\ x_{2}\\ \vdots\\ x_{m}\end{array}\right)$ (4)
$=\left(\begin{array}[]{c}w_{1}x_{1}+w_{21}x_{2}+\cdots w_{m,1}x_{m}\\ w_{1,2}x_{1}+w_{2,2}x_{2}+\cdots w_{m,2}x_{m}\\ \vdots\\ w_{1,n}x_{1}+w_{2,n}x_{2}+\cdots w_{mn}x_{m}\end{array}\right)$
Let $\mathbf{W}\in\mathbb{R}^{(m,n)}$ and $\mathbf{x}\in\mathbb{R}^{(m,1)}$, the Diamond Matrix Multiplication is defined as $\mathbf{y}=\mathbf{W}\diamond\mathbf{x}$, where $\mathbf{y}\in\mathbb{R}^{(m,1)}$. The computation procedure involves element-wise multiplication and summation of corresponding elements from left to right across the columns of matrix $\mathbf{W}$ with vector $\mathbf{x}$. The sum of the multiplication of the $i$th column of $\mathbf{W}$ with $\mathbf{x}$ is assigned as the $i$th element of $\mathbf{y}$. A detailed calculation process is provided in Eq. (4). The computation process for $\mathbf{y}=\mathbf{x}\diamond\mathbf{W}$ is identical to that of $\mathbf{W}\diamond\mathbf{x}$, implying that $\mathbf{W}\diamond\mathbf{x}=\mathbf{x}\diamond\mathbf{W}$.
Furthermore, based on its computation method, it can be inferred that the Diamond Matrix Multiplication is related to conventional matrix multiplication, specifically: $\mathbf{W}\diamond\mathbf{x}=\mathbf{W}^{T}\mathbf{x}$. Additionally, when $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$ are square matrices, we can derive the following relationships: $\mathbf{W}_{1}\diamond[\mathbf{x}\diamond\mathbf{W}_{2}]=\mathbf{W}_{2}^{T}\diamond\mathbf{W}_{1}\diamond\mathbf{x}$ and $[\mathbf{W}_{1}\diamond\mathbf{x}]\diamond\mathbf{W}_{2}=\mathbf{W}_{1}^{T}\diamond\mathbf{W}_{2}\diamond\mathbf{x}$. Detailed derivations can be found in Appendix A.
3.3 Matrix-Vector Method for Linear
Figure 2: The process of converting a linear transformation into its corresponding matrix-vector representation. a: Depicts the general form of a linear transformation. b: Presents a straightforward example of a linear transformation. c: Demonstrates the transformation of the linear operation from b into the matrix-vector format.
In this section, we present a way to transform Linear operation into matrix-vector format. Figure 2 illustrates this process: Figure 2.a shows the linear transformation of multi-channel input: $\mathbf{W}\mathbf{x}=\mathbf{y}$. Figure 2.b provides a specific example of Figure 2.a, Figure 2.c converts the linear transformation in Figure 2.b into the corresponding matrix-vector representation: $\mathbf{W}^{\prime}\diamond\mathbf{x}^{\prime}=\mathbf{y}^{\prime}$. Thus, the linear transformation can be represented in matrix-vector form as follows:
$\mathbf{x}_{i+1}=\mathbf{W}_{i}\mathbf{x}^{i}\rightarrow\mathbf{x}_{i+1}^{\prime}$ $=\mathbf{W}_{i}^{\prime}\diamond\mathbf{x}_{i}^{\prime}=(\mathbf{W}_{i}^{\prime})^{T}\mathbf{x}_{i}^{\prime}$ (5)
Here, $\mathbf{x}_{i}\in\mathbb{R}^{(N,M)}$ and $\mathbf{x}_{i+1}\in\mathbb{R}^{(N,M)}$ represent the input and output of layer $i$, respectively, while $\mathbf{W}_{i}\in\mathbb{R}^{(N,N)}$ represents the parameters of layer $i$. $\mathbf{x}_{i}^{\prime}$, $\mathbf{x}_{i+1}^{\prime}$, and $\mathbf{W}_{i}^{\prime}$ are generated based on $\mathbf{x}^{i}$, $\mathbf{x}^{i+1}$, and $\mathbf{W}_{i}$ using the Matrix-Vector Method. For convenience, we use $\mathbf{W}_{i}^{\prime}$ to represent $(\mathbf{W}_{i}^{\prime})^{T}$. Therefore, $\mathbf{x}_{i+1}^{\prime}=\mathbf{W}_{i}^{\prime}\mathbf{x}_{i}^{\prime}$. So the FFN in the Transformer could be written as:
$FFN(\mathbf{x})$ $=\mathbf{W}_{2}\sigma(\mathbf{W}_{1}\mathbf{x}+\mathbf{b}_{1})+\mathbf{b}_{2}$ (6)
$\rightarrow\mathbf{W}_{2}^{\prime}\sigma(\mathbf{W}_{1}^{\prime}\mathbf{x}^{\prime}+\mathbf{b}_{1}^{\prime})+\mathbf{b}_{2}^{\prime}$
Figure 3: The Transformer process.
Figure 4: The process of transforming $\hat{\mathbf{H}}=\text{Concat}(\hat{\mathbf{H}}_{1}...\hat{\mathbf{H}}_{8})$ in the MHA into its corresponding matrix-vector form $\mathbf{W}_{HV}^{\prime}\diamond\mathbf{x}^{\prime}=\hat{\mathbf{H}}^{\prime}$.
Figure 5: This diagram illustrates the process of transforming $\text{Concat}(\hat{\mathbf{H}}_{1}...\hat{\mathbf{H}})\mathbf{W}_{O}$ in the MHA into its corresponding matrix-vector form $\mathbf{W}_{HVO}^{\prime}\diamond\mathbf{x}^{\prime}=\mathring{\mathbf{H}}^{\prime}$, different colors in the diagram correspond to different variables.
3.4 Matrix-Vector Method for MHA
We now employ the Matrix-Vector Method to elucidate the inner workings of the MHA. The mechanism is defined by the following equation:
$\mathring{\mathbf{H}}=\operatorname{MultiHead}(\mathbf{Q},\mathbf{K},\mathbf{V})$ (7)
$=$ $\operatorname{Concat}\left(\hat{\mathbf{H}}_{1},\ldots,\hat{\mathbf{H}}_{h}\right)\mathbf{W}_{O}=\hat{\mathbf{H}}\mathbf{W}_{O}$
$\operatorname{Attention}(\mathbf{x}_{i}\mathbf{W}_{Qi}=\mathbf{Q}_{i},\mathbf{x}_{i}\mathbf{W}_{Ki}=\mathbf{K}_{i},$ (8)
$\mathbf{x}_{i}\mathbf{W}_{Vi}=\mathbf{V}_{i})=\operatorname{softmax}\left(\frac{\mathbf{Q}_{i}\mathbf{K}_{i}^{T}}{\sqrt{M}}\right)\mathbf{V}_{i}$
$=\mathbf{H}_{i}\mathbf{V}_{i}=\mathbf{H}_{i}[\mathbf{x}_{i}\mathbf{W}_{Vi}]=\hat{\mathbf{H}}_{i}$
Here, $h$ represents the number of attention heads, and the input $\mathbf{x}\in\mathbb{R}^{(N,M)}$ is divided into $\mathbf{x}_{1},...,\mathbf{x}_{h}$ based on $h$. The parameters $\mathbf{W}_{Qi},\mathbf{W}_{Ki}$, and $\mathbf{W}_{Vi}$ correspond to $\mathbf{x}_{i}$. The whole process of MHA can be represented in Figure 3. Figure 3.a represents that the input $\mathbf{x}$ is split into $\mathbf{x}_{1}\cdots\mathbf{x}_{8}$ based on the number of heads. Figure 3.b represents $\mathbf{x}_{i}\mathbf{W}_{Qi}=\mathbf{Q}_{i},\mathbf{x}_{i}\mathbf{W}_{Ki}=\mathbf{K}_{i},\mathbf{x}_{i}\mathbf{W}_{Vi}=\mathbf{V}_{i}$. Figure 3c, d, e represent the process of $\operatorname{softmax}\left(\frac{\mathbf{Q}_{i}\mathbf{K}_{i}^{T}}{\sqrt{M}}\right)\mathbf{V}_{i}$ and we omit $\sqrt{M}$. Figure 3.f represents $\operatorname{Concat}\left(\hat{\mathbf{H}}_{1},\ldots,\hat{\mathbf{H}}_{h}\right)\mathbf{W}_{O}$. Figure 3.g represents the whole process by a matrix multiplication of $(\mathbf{W}^{\prime}_{HVO})^{T}\mathbf{x}^{\prime}=\mathring{\mathbf{H}}^{\prime}$, where $\mathbf{W}^{\prime}_{HVO}$ is generated based $\mathbf{H}_{1}\cdots\mathbf{H}_{8}$, $\mathbf{W}_{V1}\cdots\mathbf{W}_{V8}$ and $\mathbf{W}_{O}$. This means that we could represent the whole complex MHA in matrix multiplication. Next, we give the proof of this process.
Our objective is to express the MHA as $(\mathbf{W}^{\prime}_{HVO})^{T}\mathbf{x}^{\prime}=\mathring{\mathbf{H}}^{\prime}$. Prior to transforming the MHA into the matrix-vector form, we need to conduct a comprehensive analysis and clearly define the research object. In Eq. (7), $\hat{\mathbf{H}}=\operatorname{Concat}\left(\hat{\mathbf{H}}_{1},\ldots,\hat{\mathbf{H}}_{h}\right)$ describes an engineering process that requires mathematical representation. The learning process for the input $\mathbf{x}$ primarily consists of two parts: $\mathbf{V}_{1},\ldots,\mathbf{V}_{h}$ and $\mathbf{H}_{1},\ldots,\mathbf{H}_{h}$. $\mathbf{H}_{i}$ is derived based on the parameters $\mathbf{x}_{i}$, $\mathbf{W}_{Qi}$ and $\mathbf{W}_{Ki}$.
Figure 4.a represents the computation process of $\mathbf{H}_{1}\mathbf{V}_{1}=\hat{\mathbf{H}}_{1},\ldots,\mathbf{H}_{h}\mathbf{V}_{h}=\hat{\mathbf{H}}_{h}$. In Figure 4.b, we present a simple example of Figure 4.a. Figure 4.b depicts $\mathbf{H}_{i}[\mathbf{x}_{i}\mathbf{W}_{Vi}],i=1,2$. In Figure 4.c we convert Figure 4.b into the matrix-vector form $\mathbf{W}_{HV}^{\prime}\diamond\mathbf{x}^{\prime}=\hat{\mathbf{H}}^{\prime}$, where $\mathbf{W}_{HV}^{\prime}$ is generated from $\mathbf{H}_{1},\mathbf{H}_{2}$ and $\mathbf{W}_{V1},\mathbf{W}_{V2}$. More details can be found in Appendix B.
Figure 5 illustrates the parameter transformation scenario after incorporating $\mathbf{W}_{O}$. Figure 5.a demonstrates an example of $\mathring{\mathbf{H}}=\operatorname{Concat}\left(\hat{\mathbf{H}}_{1},\ldots,\hat{\mathbf{H}}_{h}\right)\mathbf{W}_{O}$, while Figure 5.b rewrites Figure 5.a as $\hat{\mathbf{H}}^{\prime}\diamond\mathbf{W}_{O}^{\prime}=\mathring{\mathbf{H}}^{\prime}$. Consequently, based on Figures 4 and 5, the entire MHA can be expressed as $\mathbf{W}_{HVO}^{\prime}\diamond\mathbf{x}^{\prime}=\mathring{\mathbf{H}}^{\prime}$, where $\mathbf{W}_{HVO}^{\prime}=\mathbf{W}_{HV}^{T}\diamond\mathbf{W}_{O}^{\prime}$ (Based on Figures 5, we can get $\mathbf{W}_{HVO}^{\prime}$, but the resulting expression is too complex and not easy to understand, so we did not compute the final result. However, it is easy to analyze that as long as the variables in $\mathbf{W}_{HV}^{T}$ and $\mathbf{W}_{O}^{\prime}$ are non-zero, the result is essentially a dense matrix). So the whole MHA can be written as:
$\mathring{\mathbf{H}}^{\prime}$ $=(\mathbf{W}_{HVO}^{\prime})^{T}\mathbf{x}^{\prime}$ (9)
Here, $\mathbf{x}\in\mathbb{R}^{(NM,1)}$ and $\mathring{\mathbf{H}}\in\mathbb{R}^{(NM,1)}$ are the input and output, respectively, while $\mathbf{W}_{HVO}^{\prime}\in\mathbb{R}^{(NM,NM)}$ denotes matrices generated in accordance with $\mathbf{H}$, $\mathbf{W}_{V}$ and $\mathbf{W}_{O}$. In this manner, we have expressed MHA as a matrix-vector multiplication. This matrix multiplication representation provides a more concise way to express the MHA mechanism.
For convenience, we use $\mathbf{W}^{\prime}$ to represent $(\mathbf{W}_{HVO}^{\prime})^{T}$. The MHA operation could be simply written as:
$\mathring{\mathbf{H}}^{\prime}=\mathbf{W}^{\prime}\mathbf{x}^{\prime}$ (10)
3.5 The UAT Format of Multi-Layer Transformer
The Transformer architecture is founded on two pivotal components: FFN and MHA. In Sections 3.3 and 3.4, we have showcased the matrix-vector representations for both FFN and MHA. In this section, we delve into why multi-layer Transformer is the implementation of UAT.
Based on Equations 6 and 10, we can demonstrate that $i+1$-th layer ($i=0,1,2,\ldots$) Transformer is a concrete realization of the UAT, which can be written as (see Appendix C for more details):
$\mathbf{x}_{i+1}=(\mathbf{W}_{i+1,1}^{\prime}\mathbf{x}_{0}+\mathbf{b}_{i+1,1})+\sum_{j=1}^{i+1}\mathbf{W}^{\prime}_{j,3}\sigma(\mathbf{W}^{\prime}_{j,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{j,2})$ (11)
where $\mathbf{x}_{i+1}$ is the output of ${i+1}$-th layer, $\mathbf{x}_{0}$ is the input of the network, $\mathbf{b}^{\prime}_{j,2}=(\mathbf{W}^{\prime}_{j,2}\mathbf{b}^{\prime}_{j-1,3}+\mathbf{b}^{\prime}_{j,2})+\mathbf{W}^{\prime}_{j,2}UAT_{j-1}^{R}$, where $UAT_{j-1}^{R}=\sum_{k=1}^{j-1}\mathbf{W}^{\prime}_{k,3}\sigma(\mathbf{W}^{\prime}_{k,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{k,2})$. The term $\mathbf{b}^{\prime}_{j,2}$ is approximated by the $j$ layer of UAT with $\mathbf{x}_{0}$ as input. This enhances the model’s ability to dynamically adjust functions based on input. In the MHA mechanism, the parameters change dynamically with the input. Therefore, in the formula above, all $\mathbf{W}^{\prime}_{j,1}$, $\mathbf{W}^{\prime}_{j,2}$, and $\mathbf{W}^{\prime}_{j,3}$ parameters in layer $i$, where $j=1,\ldots,i$, are dynamically adjusted based on the input.
4 Discussion
Leveraging on our proof in the previous section that the Transformer is the tangible embodiment of UAT, in this section we address the following critical problems and explore the technical strategies of LLMs: Why does the Transformer architecture possess such power in enabling intelligent language models, such as translation and programming? What enables LLMs’ capacity for ICL? How does the LoRA scheme effectively fine-tune LLMs? What justifies the feasibility of pruning LLMs?
4.1 What Makes Transformer So Powerful in LLMs?
Theoretically, we have established that Transformer networks are concrete implementations of the UAT, enabling them to approximate any Borel measurable function. While UAT provides powerful function approximation capabilities, it inherently lacks the ability to approximate multiple functions simultaneously. However, language tasks are inherently diverse, often requiring the approximation of different functions based on the input. For instance, when summarizing, translating, or continuing the same text, the input functions are nearly identical, with only minor variations in the prompt. Without the ability to dynamically approximate functions based on the input, simply fitting a general function trend based on input will result in identical or similar outputs.
To address this, LLMs must distinguish and adapt to these nearly identical functions, dynamically generating response functions based on the input. The MHA and residual mechanisms in Transformers equip LLMs with the ability to dynamically approximate relevant functions according to the input. Specifically, MHA allows for the dynamic adjustment of the weight parameters in UAT in response to the input, while the residual mechanism supports the dynamic approximation of bias terms, as illustrated in Figure 6. This theoretical foundation enables Transformer-based LLMs to handle a wide range of tasks, including translation, continuation, summarization, code generation, and solving mathematical problems.
Figure 6: This diagram vividly demonstrates the output differences between an MLP and a Transformer when given the same input content (represented by the blue line in a) with only slight variations (represented by the black line in a). Since the input content is nearly identical and the MLP’s parameters are fixed, the MLP tends to produce the same output (b), influenced by the overall input function trend. However, the Transformer, with its ability to dynamically adapt to the input, generates different outputs based on the input variations. In the context of LLMs, the blue line in a can be understood as a consistent piece of text, while the black line represents different prompts, such as translation, continuation, or summarization.
Furthermore, the MHA mechanism can capture global information (as shown in Figure 5, where each element in the output $\mathring{\mathbf{H}}^{\prime}$ contains global information), which helps generate content that is consistent with the context. This is crucial for language understanding that requires extensive contextual information. In contrast, 1D convolution employs a sliding convolution learning strategy, where the learned content is somewhat influenced by the size of the convolution kernel, leading to a more fragmented learning approach (for more information, please refer to Appendix D).
4.2 What Enables LLMs to Possess ICL Capability?
Contextual interaction, as the core capability of LLMs, permeates every phase from training and fine-tuning to prediction. ICL, multi-step reasoning, and instruction following are intuitive manifestations of this contextual interaction. Leveraging their context-sensitive interaction capabilities, LLMs can exhibit behaviors consistent with ICL, multi-step inference, and instruction following, which are tailored based on contextual cues.
So, how does this contextual interaction capability arise within LLMs? The formula $\mathring{\mathbf{H}}=\mathbf{W}^{\prime}\mathbf{x}^{\prime}$ in Figure 5 reveals this mode of contextual interaction. Since $\mathbf{W}^{\prime}$ represents a dense matrix (almost devoid of zero elements and whose internal elements are highly correlated), each element in $\mathring{\mathbf{H}}$ encapsulates comprehensive information from both preceding and subsequent contexts. This learning of holistic contextual information constitutes the foundation of contextual interaction within LLMs. (See Appendix D for more details)
4.3 What Justifies the Feasibility of Pruning LLMs?
Due to the massive size of parameters in LLMs and the subsequent high demand for computational resources, pruning LLMs is pivotal for their deployment. A legitimate question to ask is why LLMs are amenable to pruning. The rationale lies in the presence of excessively low-weight parameters in certain layers of LLMs. To understand this, we can directly analyze it from the perspective of the formula underlying the UAT:
$|\sum_{j=1}^{N}\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)-f(\mathbf{x})|<\varepsilon\quad.$ (12)
for all $\mathbf{x}\in\mathbf{I}_{n}$. Let’s assume $\Lambda=\{1,2...N\},\Lambda_{1}\cup\Lambda_{2}=\Lambda,\Lambda_{1}\cap\Lambda_{2}=\emptyset$ and $|\sum_{j\in\Lambda_{1}}\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)|\to 0$. Then we have:
$|\sum_{j\in\Lambda_{1}}\alpha_{j}\sigma\left(\mathbf{W}j^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)$ (13)
$+$ $\sum_{j\in\Lambda_{2}}\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)-f(\mathbf{x})|<\varepsilon$
Since $|\sum_{j\in\Lambda_{1}}\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)|\to 0$, we have the following inequality:
$\big{|}|\sum_{j\in\Lambda_{1}}\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)-f(\mathbf{x})\mid$ (14)
-- $|\sum_{j\in\Lambda_{2}}\alpha_{j}\sigma\left(\mathbf{W}j^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)|\big{|}<\varepsilon$
Therefore, we have:
$|\sum_{j=1}^{N_{1}}\alpha_{j}\sigma\left(\mathbf{W}j^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)-f(\mathbf{x})|$ (15)
$<\varepsilon+|\sum_{j=1}^{N_{2}}\alpha_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\theta_{j}\right)|$
Hence, when parameters in certain layers are small enough, we can directly remove those layers since their impact on the final result is minimal.
4.4 How Does the LoRA Scheme Effectively Fine-tune LLMs?
Given the substantial computational resources required to train LLMs and their powerful generalization abilities, we believe that more efficient use of pre-trained models is essential. Re-training models from scratch incurs significant computational costs, so reusing well-trained models for new tasks is both practical and resource-efficient. A prominent solution to this challenge is the LoRA (Hu et al. 2021), which can be expressed as follows:
$\mathbf{h}=\mathbf{W}_{0}\mathbf{x}+\Delta\mathbf{W}\mathbf{x}$ (16)
According to Eq. 11, we use LoRA to fine-tune the Linear operation in FFN can be written as:
$\mathbf{x}_{i+1}$ $=(\mathbf{W}_{i+1,1}^{\prime}\mathbf{x}_{0}+\mathbf{b}_{i+1,1})+\sum_{j=1}^{i+1}(\mathbf{W}^{\prime}_{j,3}+\Delta\mathbf{W}_{j+1,3})$ (17)
$\sigma((\mathbf{W}^{\prime}_{j,2}+\Delta\mathbf{W}_{j+1,2})\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{j,2})$
From Eq. 17, it can be seen that LoRA essentially fine-tunes the UAT parameters layer by layer for a specific task.
Figure 7: The comparison between Human and LLMs.
4.5 Rethinking LLMs
The capabilities of LLMs have become so advanced that their language processing abilities are approaching human levels, raising a core question: How do LLMs differ from humans in language processing? Figure 7 illustrates the comparison between the language processing processes of LLMs and humans. Both start with language encoding—humans encode language through a character-based system, while LLMs use numerical arrays. At this level, there is almost no difference. Given the ambiguity of words, determining context is crucial: humans understand context through the activation and transmission of neurons in the brain, while LLMs approximate the corresponding functions using UAT. Here, the input and output of network layers are analogous to the transmission of neural signals in the brain, and the function approximation corresponds to the final output of humans. From this perspective, the differences between humans and LLMs in language processing seem minimal.
So how do we explain human understanding, analysis of language, and memory retrieval? Are these also capabilities of LLMs? First, we can consider the human brain as a combination of one or more UAT models, which are randomly initialized at birth. What supports such an assumption? The original neural network, the perceptron (Rosenblatt 1963), was designed based on the human neuron. Over time, as deep learning networks evolved, the difference in form between deep networks and the perceptron grew, leading people to no longer associate neural networks with human neurons. However, through our derivation, the Transformer can also be understood as an implementation of UAT, and the perceptron can be seen as a layer of UAT. The human brain can be considered as a cluster of multi-layer perception. Learning language in a social environment is akin to training the UATs in the brain. Therefore, what we call understanding, analysis, and memory retrieval are essentially processes of fitting outputs based on inputs. Understanding and analysis: When we respond to a question posed by someone, it can be seen as a simple example of understanding and analysis, which both humans and LLMs can accomplish. Humans may guess the answer to unfamiliar questions based on their experiences, which can be either correct or incorrect, and LLMs exhibit similar behavior. Memory retrieval: This is essentially fitting specific results based on particular words, as there is no actual database in the brain. For instance, recalling the experience of eating an apple for the first time in childhood is the brain fitting specific results based on those words; without those words, the memory would not surface. This is because the brain dynamically outputs results based on inputs and fills in the details of the event based on learned natural rules (We provide some examples about understanding, analysis of language, and memory retrieval in Appendix E). However, as humans grow, the weights in the brain are constantly updated, leading to potential memory distortions.
Therefore, it is entirely reasonable for LLMs to make errors or generate hallucinations—these are just outputs produced based on existing weights and inputs, a problem humans also face. We believe the greatest advantage humans have over LLMs is their powerful multimodal and multitask processing abilities (which can be understood as the coordination of multiple UAT models to produce reasonable results). Another advantage is the interaction with the real world, which allows us to verify the knowledge we’ve learned in reality, enabling the brain to continuously optimize its parameters based on inputs. LLMs, on the other hand, are limited to function approximation within the corpora data.
5 Conclusion
In this paper, we delve into the theoretical underpinnings of LLMs, demonstrating that contemporary LLMs, primarily constructed with Transformer architectures, embody concrete manifestations of the UAT. The remarkable generalization prowess exhibited by LLMs is attributed to their MHA modules and residual operation, which enable the adaptation to approximate diverse functions based on the presented input data. Contextual interaction emerges as a paramount capability for LLMs, manifesting such abilities as ICL, instruction following, and contextual reasoning. These competencies are enabled by the Transformer’s MHA to learn from context.
Expanding upon this understanding, we have provided a rigorous theoretical grounding for key techniques employed in LLMs, including LoRA for efficient fine-tuning and pruning for model compression, elucidating their effectiveness through the lens of the UAT. By leveraging the theoretical framework provided by the UAT, not only can existing methodologies be explained but also avenues for the future evolution of LLMs are illuminated.
Appendix A The Property of Diamond
[…]
$\mathbf{W}_{1}\diamond[\mathbf{x}\diamond\mathbf{W}_{2}]$ (1)
$=$ $\mathbf{W}_{1}\diamond[\mathbf{W}_{2}^{T}\mathbf{x}]$
$=$ $\mathbf{W}_{1}^{T}[\mathbf{W}_{2}^{T}\mathbf{x}]$
$=$ $\mathbf{W}_{1}^{T}\mathbf{W}_{2}^{T}\mathbf{x}=\mathbf{W}_{12}\mathbf{x}$
$=$ $(\mathbf{W}_{1}^{T}\mathbf{W}_{2}^{T})^{T}\diamond\mathbf{x}$
$=$ $(\mathbf{W}_{2}\mathbf{W}_{1})\diamond\mathbf{x}$
$=$ $\mathbf{W}_{2}^{T}\diamond\mathbf{W}_{1}\diamond\mathbf{x}$
$[\mathbf{W}_{1}\diamond\mathbf{x}]\diamond\mathbf{W}_{2}$ (2)
$=$ $[\mathbf{W}_{1}^{T}\mathbf{x}]\diamond\mathbf{W}_{2}$
$=$ $\mathbf{W}_{2}^{T}[\mathbf{W}_{1}^{T}\mathbf{x}]$
$=$ $[\mathbf{W}_{2}^{T}\mathbf{W}_{1}^{T}]^{T}\diamond\mathbf{x}$
$=$ $[\mathbf{W}_{1}\mathbf{W}_{2}]\diamond\mathbf{x}$
$=$ $\mathbf{W}_{1}^{T}\diamond\mathbf{W}_{2}\diamond\mathbf{x}$
Figure 1: The process of transforming $\text{Concat}(\hat{\mathbf{H}}_{1}...\hat{\mathbf{H}}_{8})$ in the MHA into its corresponding matrix-vector form $\mathbf{W}_{HV}^{\prime}\mathbf{x}^{\prime}=\hat{\mathbf{H}}^{\prime}$.
Figure 2: The process of Transformer.
Appendix B The Matrix-Vector Form of MHA
In this section, we will present the transformation process of MHA into their corresponding matrix-vector forms.
Figure 1.a represents the computation process of obtaining $\mathbf{W}_{QKi}$ in the attention mechanism, while Figure 1.b computes $\mathbf{H}_{1},\dots,\mathbf{H}_{8}$ through the softmax operation. Figure 1.c and Figure 1.e illustrates the process of converting $\operatorname{Concat}\left(\mathbf{H}_{1}\mathbf{V}_{1},\ldots,\mathbf{H}_{h}\mathbf{V}_{h}\right)$ into the matrix-vector form $\mathbf{W}_{HV}^{\prime}\diamond\mathbf{x}^{\prime}=\hat{\mathbf{H}}^{\prime}$, where $\mathbf{W}_{HV}^{\prime}$ is generated from $\mathbf{H}_{1},\dots,\mathbf{H}_{8}$ and $\mathbf{W}_{V1},\dots,\mathbf{W}_{V8}$. Figure 1.c shows $\mathbf{H}_{i}[\mathbf{x}_{i}\mathbf{W}_{Vi}]$. In Figure 1.d, we provide a simple example demonstrating the conversion of $\mathbf{H}_{i}[\mathbf{x}_{i}\mathbf{W}_{Vi}]$ into $\mathbf{W}_{HVi}^{\prime}\mathbf{x}_{i}^{\prime}$. This process is divided into four parts: Figure 1.d.1 represents the general form of $\mathbf{H}_{i}[\mathbf{x}_{i}\mathbf{W}_{Vi}]$, while Figure 1.d.2 serves as a simple example of Figure 1.d.1. Figure 1.d.3 first rewrites Figure 1.d.2 using diamond multiplication as $\mathbf{H}_{i}^{\prime}\diamond[\mathbf{x}_{i}^{\prime}\diamond\mathbf{W}_{Vi}^{\prime}]$, and then utilizes the property of diamond multiplication to express it as $(\mathbf{W}_{Vi}^{\prime})^{T}\diamond\mathbf{H}_{i}^{\prime}\diamond\mathbf{x}_{i}^{\prime}$ in Figure 1.d.4. Finally, Figure 1.d.5 is obtained as $\mathbf{W}_{HVi}^{\prime}\diamond\mathbf{x}_{i}^{\prime}=\mathbf{H}_{i}^{\prime}$. It can be observed that $\mathbf{W}_{HVi}^{\prime}$ is a dense matrix. In Figure 1.e, we present a simple example of $\mathbf{H}_{1}[\mathbf{x}_{1}\mathbf{W}_{V1}],\mathbf{H}_{2}[\mathbf{x}_{2}\mathbf{W}_{V2}]...$ into the matrix-vector form $\mathbf{W}_{HV}^{\prime}\mathbf{x}^{\prime}$. Figure 1.e.1 depicts $\mathbf{H}_{i}[\mathbf{x}_{i}\mathbf{W}_{Vi}]$, while Figure 1.e.2 represents the Matrix-vector format of Figure 1.e.1.
Appendix C The UAT Format of Transformer
In this section, we will prove multi-layer Transformer is the implementation of UAT.
C.1 The Properties of UAT
Before expressing Transformers in the UAT format, we present two lemma regarding UAT. There are two cases for UAT-approximated functions: $f(\mathbf{x})\in\mathbb{R}$ and $f(\mathbf{x})\in\mathbb{R}^{m}$. The proof for the case where $f(\mathbf{x})\in\mathbb{R}$ can be inferred from $f(\mathbf{x})\in\mathbb{R}^{m}$. Therefore, we will only provide the proof for approximating $f(\mathbf{x})\in\mathbb{R}^{m}$ using UAT.
Lemma 1. The mathematical form of UAT remains unchanged when multiplied by a matrix (constant).
$G(\mathbf{x})$ $=\mathbf{\beta}\sum_{j=1}^{N}\mathbf{\alpha}_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\mathbf{\theta}_{j}\right)$ (3)
$=\sum_{j=1}^{N}\mathbf{\beta}\mathbf{\alpha}_{j}\sigma\left(\mathbf{W}_{j}^{\mathrm{T}}\mathbf{x}+\mathbf{\theta}_{j}\right)$
Eq. (3) shows the representation of UAT multiplying a matrix. Let $\mathbf{\alpha}_{j}=\mathbf{\beta}\mathbf{\alpha}_{j}$, and the result remains consistent with the original UAT mathematical form. Thus, it is proven that the mathematical form of UAT remains unchanged when multiplied by a matrix (constant).
Lemma 2. Let the general term of the network be written as: $\mathbf{x}^{\prime}_{i+1}=(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})+\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$, $i=0,1,2,3,\ldots$. The multi-layer network of this form still corresponds to the mathematical form of UAT, and we refer to such a network as a residual network.
To prove that a multi-layer network with the general term written as $\mathbf{x}^{\prime}_{i+1}=(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})+\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$ corresponds to the UAT mathematical form, we first provide the forms of single-layer and two-layer networks, as shown in Eq. (4) and (5).
$\mathbf{x}^{\prime}_{1}=(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,3})+\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})$ (4)
$\mathbf{x}^{\prime}_{2}$ $=\mathbf{W}^{\prime}_{2,1}\mathbf{x}^{\prime}_{1}+\mathbf{W}^{\prime}_{2,3}\sigma(\mathbf{W}^{\prime}_{2,2}\mathbf{x}^{\prime}_{1}+\mathbf{b}^{\prime}_{2,2})+\mathbf{b}^{\prime}_{2,3}$ (5)
$=(\mathbf{W}^{\prime}_{2,1}\mathbf{x}^{\prime}_{1}+\mathbf{b}^{\prime}_{2,3})+\mathbf{W}^{\prime}_{2,3}\sigma(\mathbf{W}^{\prime}_{2,2}\mathbf{x}^{\prime}_{1}+\mathbf{b}^{\prime}_{2,2})$
=$=\{\mathbf{W}^{\prime}_{2,1}[(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,3})+\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})]+\mathbf{b}^{\prime}_{2,3}\}$
$+\mathbf{W}^{\prime}_{2,3}\sigma\{\mathbf{W}^{\prime}_{2,2}[(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,3})+\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})]+\mathbf{b}^{\prime}_{2,2}\}$
=$=\{\mathbf{W}^{\prime}_{2,1}(\mathbf{W}^{\prime}_{1,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,3})+\mathbf{b}^{\prime}_{2,3}+\mathbf{W}^{\prime}_{2,1}\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})\}$
$+\mathbf{W}^{\prime}_{2,3}\sigma\{\mathbf{W}^{\prime}_{2,2}({\mathbf{W}}^{\prime}_{1,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,3})+\mathbf{b}^{\prime}_{2,2}+\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})\}$
=$=\{(\underline{\mathbf{W}^{\prime}_{2,1}\mathbf{W}^{\prime}_{1,1}}\mathbf{x}^{\prime}_{0}+\underline{\mathbf{W}^{\prime}_{2,1}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,3}})+\underline{\mathbf{W}^{\prime}_{2,1}\mathbf{W}^{\prime}_{1,3}}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})\}$
$+\mathbf{W}^{\prime}_{2,3}\sigma\{\underline{\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,1}}\mathbf{x}^{\prime}_{0}+\underline{(\mathbf{W}^{\prime}_{2,2}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,2})+\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})}\}$
In Eq. (5), let $\mathbf{W}^{\prime}_{2,1}=\underline{\mathbf{W}^{\prime}_{2,1}\mathbf{W}^{\prime}_{1,1}}$, $\mathbf{b}^{\prime}_{2,3}=\underline{\mathbf{W}^{\prime}_{2,1}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,3}}$, $\mathbf{W}^{\prime}_{1,3}=\underline{\mathbf{W}^{\prime}_{2,1}\mathbf{W}^{\prime}_{1,3}}$, $\mathbf{W}^{\prime}_{2,3}=\underline{\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,1}}$, and $\mathbf{b}^{\prime}_{2,2}=\underline{(\mathbf{W}^{\prime}_{2,2}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,2})+\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})}$. Thus, Eq. (6) can be written as:
$\mathbf{x}^{\prime}_{2}$ $=(\mathbf{W}^{\prime}_{2,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{2,1})+\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})$ (6)
$+\mathbf{W}^{\prime}_{2,3}\sigma(\mathbf{W}^{\prime}_{2,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{2,2})$
Why can we combine the above terms? Once the network training is complete, we know these parameters, so they can be directly calculated and thus combined. And $\mathbf{b}^{\prime}_{2,2}=\underline{(\mathbf{W}^{\prime}_{2,2}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,2})+\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})}$ clearly has the same mathematical form as UAT. We can understand it as dynamically fitting the bias term $\mathbf{b}^{\prime}_{2,2}$ using UAT based on the input. Therefore, we have proven that the mathematical form of one and two-layer residual networks is consistent with the UAT mathematical form.
Assume that the mathematical form of the first $i$ layers of the residual network is consistent with UAT. Our goal is to prove that the mathematical form of the $i+1$-th layer of the residual network is still consistent with UAT. For the convenience of expression, we make the following definition: since the mathematical form of the first $i$ layers of the residual network is consistent with UAT, we write $\mathbf{x}^{\prime}_{i}$ as $\mathbf{x}^{\prime}_{i}=(\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{i,3})+UAT^{R}_{i}$, where the first term is written separately and the rest is written as the remainder term $UAT^{R}_{i}$. Since $\mathbf{x}^{\prime}_{i+1}=(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})+\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$, we divide $\mathbf{x}^{\prime}_{i+1}$ into two parts: $(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})$ and $\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$.
First, consider the part $(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})$. Substitute $\mathbf{x}^{\prime}_{i}=(\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{i,3})+UAT^{R}_{i}$ into it, and we get Eq. (7). Let $\mathbf{W}^{\prime}_{i+1,1}=\underline{\mathbf{W}^{\prime}_{i+1,1}\mathbf{W}^{\prime}_{i,1}}$ and $\mathbf{b}^{\prime}_{i+1,3}=\underline{\mathbf{W}^{\prime}_{i+1,1}\mathbf{b}^{\prime}_{i,1}+\mathbf{b}^{\prime}_{i+1,3}}$, we can simplify the first part to $(\mathbf{W}^{\prime}_{i+1,}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{i+1,3})+\mathbf{W}^{\prime}_{i+1,1}UAT^{R}_{i}$. Since the mathematical form of $UAT^{R}_{i}$ is consistent with UAT, and we have already proven that the mathematical form of UAT remains unchanged when multiplied by a matrix, we have proven that the mathematical form of the first part is consistent with UAT.
$(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})$ (7)
$=$ $\{\mathbf{W}^{\prime}_{i+1,1}[(\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{i,1})+UAT^{R}_{i}]+\mathbf{b}^{\prime}_{i+1,3}\}$
$=$ $\{\mathbf{W}^{\prime}_{i+1,1}(\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{i,1})+\mathbf{W}^{\prime}_{i+1,1}UAT^{R}_{i}+\mathbf{b}^{\prime}_{i+1,3}\}$
$=$ $\{[\underline{\mathbf{W}^{\prime}_{i+1,1}\mathbf{W}^{\prime}_{i,1}}\mathbf{x}^{\prime}_{0}+\underline{(\mathbf{W}^{\prime}_{i+1,1}\mathbf{b}^{\prime}_{i,1}+\mathbf{b}^{\prime}_{i+1,3})}]+\mathbf{W}^{\prime}_{i+1,1}UAT^{R}_{i}\}$
Next, we prove that the mathematical form of the second part, $\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$, is consistent with UAT. By substituting $\mathbf{x}^{\prime}_{i}=(\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{i,3})+UAT^{R}_{i}$ into it, we obtain Eq. (8). Let $\mathbf{W}^{\prime}_{i+1,2}=\underline{\mathbf{W}^{\prime}_{i+1,2}\mathbf{W}^{\prime}_{i,1}}$ and $\mathbf{b}^{\prime}_{i+1,2}=\underline{(\mathbf{W}^{\prime}_{i+1,2}\mathbf{b}^{\prime}_{i,1}+\mathbf{b}^{\prime}_{i+1,2})+\mathbf{W}^{\prime}_{i+1,2}UAT^{R}_{i}}$. Thus, the second term can be written as $\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{i+1,2})$, which can be considered as a term within UAT. Here, $\mathbf{b}^{\prime}_{i+1,2}$ is a bias term approximated using UAT.
$\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$ (8)
$=$ $\mathbf{W}^{\prime}_{i+1,3}\sigma\{\mathbf{W}^{\prime}_{i+1,2}[(\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{i,1})+UAT^{R}_{i}]+\mathbf{b}^{\prime}_{i+1,2}\}$
$=$ $\mathbf{W}^{\prime}_{i+1,3}\sigma\{\mathbf{W}^{\prime}_{i+1,2}(\mathbf{W}^{\prime}_{i,1}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{i,1})+\mathbf{W}^{\prime}_{i+1,2}UAT^{R}_{i}+\mathbf{b}^{\prime}_{i+1,2}\}$
$=$ $\mathbf{W}^{\prime}_{i+1,3}\sigma\{\underline{\mathbf{W}^{\prime}_{i+1,2}\mathbf{W}^{\prime}_{i,1}}\mathbf{x}^{\prime}_{0}+\underline{(\mathbf{W}^{\prime}_{i+1,2}\mathbf{b}^{\prime}_{i,1}+\mathbf{b}^{\prime}_{i+1,2})+\mathbf{W}^{\prime}_{i+1,2}UAT^{R}_{i}}\}$
Since the first term of $\mathbf{x}^{\prime}_{i+1}$ has the same mathematical form as UAT, and the second term can be considered as a component within UAT, the sum of both terms still corresponds to the UAT mathematical form. Therefore, we have proven that the mathematical form of a multi-layer residual network is consistent with the UAT mathematical form. The difference lies in that some bias terms are approximated using UAT rather than being directly defined.
C.2 Transformer to UAT
Similarly, To illustrate relationship between multi-layer Transformer and UAT, we will first derive the general form based on Figure 2 and then compare it with the general form given in Section C.1: $\mathbf{x}^{\prime}_{i+1}=(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})+\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2}).$
By examining the relationship between these forms, we can demonstrate that the mathematical form of multi-layer Transformer networks is consistent with the UAT. The Transformer involves two key operations: MHA and FFN. These operations, when expressed in matrix-vector form, correspond to Eq.(10) and Eq.(6), respectively. Consequently, the general term for a Transformer-based network can be written as Eq. (9).
$\mathbf{x}^{\prime}_{i+1}$ $=\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{W}^{\prime}_{i+1,3}\sigma[\mathbf{W}^{\prime}_{i+1,2}(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i})+\mathbf{b}^{\prime}_{i+1,2}]+\mathbf{b}^{\prime}_{i+1,3}$ (9)
$=\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{W}^{\prime}_{i+1,3}\sigma(\underline{\mathbf{W}^{\prime}_{i+1,2}\mathbf{W}^{\prime}_{i+1,1}}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})+\mathbf{b}^{\prime}_{i+1,3}$
Let $\mathbf{W}^{\prime}_{i+1,2}=\underline{\mathbf{W}^{\prime}_{i+1,2}\mathbf{W}^{\prime}_{i+1,1}}$, so we have $\mathbf{x}^{\prime}_{i+1}=(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})+\mathbf{W}^{\prime}_{i,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$. It is evident that, when compared to the general form given in Section C.1: $\mathbf{x}^{\prime}_{i+1}=(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})+\mathbf{W}^{\prime}_{i+1,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$, the mathematical forms are the same. Therefore, multi-layer Transformers are also specific implementations of the UAT. Similarly, to clearly present the UAT format, we use the method mentioned in Section C.1. Assuming the mathematical form of $\mathbf{x}^{\prime}_{i}$ is consistent with UAT, we decompose $\mathbf{x}^{\prime}_{i}$ into a primary term plus a remainder term, written as $\mathbf{x}^{\prime}_{i}=(\mathbf{W}_{i,1}\mathbf{x}_{0}+\mathbf{b}_{i,3})+UAT^{R}_{i}$. Substituting this into $\mathbf{x}^{\prime}_{i+1}=(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})+\mathbf{W}^{\prime}_{i,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$, we obtain Eq. (11). By setting $\mathbf{W}^{\prime}_{i+1,1}=\underline{\mathbf{W}^{\prime}_{i+1,1}\mathbf{W}_{i,1}}$, $\mathbf{b}^{\prime}_{i+1,3}=\underline{(\mathbf{W}^{\prime}_{i+1,1}\mathbf{b}_{i,1}+\mathbf{b}^{\prime}_{i+1,3})}$, $UAT^{R}_{i}=\mathbf{W}^{\prime}_{i+1,1}UAT^{R}_{i}$, and $\mathbf{W}^{\prime}_{i+1,2}=\underline{\mathbf{W}^{\prime}_{i+1,2}\mathbf{W}_{i,1}}$, $\mathbf{b}^{\prime}_{i+1,2}=\underline{(\mathbf{W}^{\prime}_{i+1,2}\mathbf{b}_{i,1}+\mathbf{b}^{\prime}_{i+1,2})+\mathbf{W}^{\prime}_{i+1,2}UAT^{R}_{i}}$, we get $\mathbf{x}^{\prime}_{i+1}=(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}_{0}+\mathbf{b}^{\prime}_{i+1,3})+UAT^{R}_{i}+\mathbf{W}^{\prime}_{i,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}_{0}+\mathbf{b}^{\prime}_{i+1,2})$. Since the mathematical form of $UAT^{R}_{i}$ is consistent with multi-layer UAT, it follows that multi-layer Transformer networks are also specific implementations of UAT. In addition, $UAT^{R}_{i+1}=UAT^{R}_{i}+\mathbf{W}_{i+1,3}^{\prime}\sigma(\mathbf{W}_{i+1,2}^{\prime}\mathbf{x}_{0}+\mathbf{b}_{i+1,2})$. So, $\mathbf{x}_{i+1}=(\mathbf{W}_{i+1,1}^{\prime}\mathbf{x}_{0}+\mathbf{b}_{i+1,3})+UAT^{R}_{i+1}$, where $i=0,1,2\ldots$, $UAT_{i+1}^{R}=\sum_{j=1}^{i+1}\mathbf{W}^{\prime}_{j,3}\sigma(\mathbf{W}^{\prime}_{j,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{j,2})$, $\mathbf{b}^{\prime}_{j,2}=(\mathbf{W}^{\prime}_{j,2}\mathbf{b}^{\prime}_{j-1,3}+\mathbf{b}^{\prime}_{j,2})+\mathbf{W}^{\prime}_{j,2}UAT_{j-1}^{R}$. To sum up, the $i+1$-layer Transformer can be written as:
$\mathbf{x}_{i+1}^{\prime}=(\mathbf{W}_{i+1,1}^{\prime}\mathbf{x}_{0}^{\prime}+\mathbf{b}_{i+1,1})+\sum_{j=1}^{i+1}\mathbf{W}^{\prime}_{j,3}\sigma(\mathbf{W}^{\prime}_{j,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{j,2})$ (10)
Other parameters will also vary with the number of layers. And, these changes involve only matrix multiplications or additions. Thus, they do not fundamentally alter the mathematical form of the UAT. Therefore, we will not detail the specific forms for each layer here.
The parameters in the MHA mechanism dynamically change with the input. Therefore, in the corresponding UAT mathematical form for Transformers, all $\mathbf{W}^{\prime}_{i,1}$, $\mathbf{W}^{\prime}_{j,1}$ and $\mathbf{W}^{\prime}_{j,2}$ parameters for the $i$-th layer, where $j=1,...,i$, are dynamically changing with the input. We give some examples in Figure 3.
$\mathbf{x}^{\prime}_{i+1}$ $=(\mathbf{W}^{\prime}_{i+1,1}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,3})+\mathbf{W}^{\prime}_{i,3}\sigma(\mathbf{W}^{\prime}_{i+1,2}\mathbf{x}^{\prime}_{i}+\mathbf{b}^{\prime}_{i+1,2})$ (11)
=$=\{\mathbf{W}^{\prime}_{i+1,1}[(\mathbf{W}_{i,1}\mathbf{x}_{0}+\mathbf{b}_{i,1})+UAT^{R}_{i}]+\mathbf{b}^{\prime}_{i+1,3}\}$
$+\mathbf{W}^{\prime}_{i,3}\sigma\{\mathbf{W}^{\prime}_{i+1,2}[(\mathbf{W}_{i,1}\mathbf{x}_{0}+\mathbf{b}_{i,1})+UAT^{R}_{i}]+\mathbf{b}^{\prime}_{i+1,2}\}$
$=[\underline{\mathbf{W}^{\prime}_{i+1,1}\mathbf{W}_{i,1}}\mathbf{x}_{0}+\underline{(\mathbf{W}^{\prime}_{i+1,1}\mathbf{b}_{i,1}+\mathbf{b}^{\prime}_{i+1,3})}]+\mathbf{W}^{\prime}_{i+1,1}UAT^{R}_{i}]$
$+\mathbf{W}^{\prime}_{i,3}\sigma\{\underline{\mathbf{W}^{\prime}_{i+1,2}\mathbf{W}_{i,1}}\mathbf{x}_{0}+\underline{(\mathbf{W}^{\prime}_{i+1,2}\mathbf{b}_{i,1}+\mathbf{b}^{\prime}_{i+1,2})+\mathbf{W}^{\prime}_{i+1,2}UAT^{R}_{i}}\}$
Figure 3: Some examples of the UAT format of multi-layer Transformer. The changes of parameters are represented within the dashed boxes. The parameters on the right of the equations indicate the original values, while those on the left represent the transformed values. There is no specific order of calculation for the parameters within each dashed box, but there is a top-to-bottom calculation order between different dashed boxes.
Appendix D The Learning Process of MHA
To more clearly illustrate the learning details of the MHA mechanism in Transformers, we present an example in Figure 4. First, the input consists of 10 characters, as shown in Figure 4.a. Then, in Figure 4.b, these characters are encoded into $n$-dimensional vectors. In Figure 4.c, the encoded vectors are expanded into $\mathbf{x}^{\prime}$ using Matrix-Vector methods. Subsequently, in Figure 4.d, we compute $\mathring{\mathbf{H}}^{\prime}=(\mathbf{W}_{HVO}^{\prime})^{T}\mathbf{x}^{\prime}$. Since $(\mathbf{W}_{HVO}^{\prime})^{T}$ is an almost fully dense square matrix, each element of $\mathring{\mathbf{H}}^{\prime}$ contains information from the entire input. This is why ICL exists in LLMs: each element can access the full context.
While Figure 5 shows the learning process of 1D convolution. Compared to the holistic learning approach of MHA, 1D convolution learns by processing encoded text in chunks based on the size of the convolution kernel. This method is akin to learning the entire context each time but focusing on a partial set of characters of each word at a time.
Figure 4: The learning process of MHA
Figure 5: The learning process of 1D convolution.
Appendix E The Comparision between Human and LLMs
To more specifically compare the differences between humans and LLMs, we provide a simple example here. We prepared two questions: ”Nui zi created Nui zi dynamics” and ”Nui zi created gravitation and relativity theory.” It is important to emphasize that these sentences are entirely fictional, and there is no such person as ”Nui zi.” However, the term ”zi” has some special meanings in specific contexts that some people might know, while others may not. We have provided some explanations of ”zi” as interpreted by Chatgpt.4o in Figure 6. We have offered possible answers from different groups: a kid, a general adult, a professional researcher, and also responses from Chatgpt.4o and Qwen 2.5 to the above questions.
Figures 7 and 8 present the answers to the two questions. Both humans and LLMs provide the output based on input. To clearly indicate which works or sentences primarily affected the generated content, we used red font to label those (this part is not generated by the model). Firstly, different individuals provide different answers based on their background knowledge, which can be understood as a process of generating output based on existing training parameters in the brain and the given input. In this process, memories about keywords lead to specific outputs, such as ”Professional researcher: 1. gravitation → Isaac Newton; 2. relativity theory → Albert Einstein.” This process can be understood as understanding the question, analyzing it, retrieving memories, and producing results. However, if the human brain is considered as a UAT, it generates the output based on specific inputs, such as ”gravitation → Isaac Newton” or relevant information about the theory itself. But the truth is that the brain never creates a ’memory warehouse’ to store that information, it just produces the output based on input. So the ’memory’ could be understood as something produced by input and pretrained parameters.
Since Chatgpt.4o and Qwen 2.5 models have been trained on extensive datasets, they possess some information about ”zi,” ”Nui,” ”dynamics,” ”gravitation,” and ”relativity theory.” When these characters are inputted into the models, they use their pre-trained parameters to generate relevant outputs. For the first question, they provided possible inferences based on the keywords ”zi” and ”Nui” and highlighted the inconsistencies in ”Nui zi dynamics.” For the second question, which is more detailed, both Chatgpt.4o and Qwen 2.5 were able to provide more detailed contextual information (note that the models can generate additional content based on both the input and the output from the input), which can be understood as human-like understanding, analysis, and memory retrieval.
Figure 6: The background information about ’zi’ from Chatgpt.4o.
Figure 7: The answers about ’Nui zi created Nui zi dynamics’.
Figure 8: The answers about ’Nui zi created gravitation and relativity theory’.